Saved Help Antioch Company makes eBook readers. The company had the following amounts at the beginning of 2018: Cash, $665.000; Raw Materials Inventory, $70,000; Work in Process Inventory, $27000; Finished Goods Inventory, $58,000 Common Stock, $588,000; and Retained Earnings, $232,000. Antioch experienced the following accounting events during 2018. Other than the adjusting entries for depreciation, assume that all transactions are cash transactions. 1. Paid $28,000 of research and development costs 2. Paid $55,000 for raw materials that will be used to make eBook readers. 3. Placed $84,000 of the raw materials cost into the process of manufacturing eBook readers. 4. Paid $77,000 for salaries of selling and administrative employees 5. Paid $98,000 for wages of production workers 6. Paid $137000 to purchase equipment used in selling and administrative offices 7. Recognized depreciation on the office equipment. The equipment was acquired on January 1, 2018. It ha- d $17000 salvage value and a six-year life. The amount of depreciation is computed as [(Cost - salvage) useful life]. Specifically. ($137,000 $17,000) 6 $20,000 8. Paid $138,000 to purchase manufacturing equipment. 9. Recognized depreciation on the manufacturing equipment. The equipment was acquired on January 1, 2018. It has a $26,000 salvage value and a eight-year life. The amount of depreciation is computed as [(Cost - salvage)+ useful life). Specifically, ($138,000 $26.000) 8 $14,000. 10. Paid $47,000 for rent and utility costs on the manufacturing facility 11. Paid $79,000 for inventory holding expenses for completed eBook readers (rental of warehouse space, salaries of warehouse personnel, and other general storage cost) 12 Completed and transferred eBook readers that had total cost of $258.000 from work in process inventory to finished goods <Prev 2 3 of 3 Next > DOLL F10 F11F12 PrtScrInsert F6 F7 F9 8 9 0 5 6Hel 12 Completed and transferred eBook readers that had total cost of $258,000 from work in process inventory to finish goods. 13. Sold 830 eBook readers for $425,000 14. It cost Antioch $174,300 to make the eBook readers sold in Event 13 Problem 10-27 Part a Required a. Show how these events affect the balance sheet, income statement, and statement of cash flows by rec ling them in a horiz financial statements model. (Enter any decreases to account balances and cosh outflows with a minus sign. In the Cash Fl column, indicate whether the item is an operating activity (OA), investing activity (IA), financing activity (FA), or NA for no affected.) ANTIOCH COMPANY Financial Statements Model Stockholders Equity Assets Work in Raw ·| Process +: Goods Event Finished e ManufacturingCommon Reta No. CashMaterials Furniture Equipment Stock Inventorv 23 of 3 Next Prey DELL Priscr nsert F8 F9 F F1112 F6 0 6 7 8 5Help Save & Exit Su Check my wo 665000 70,000 + 27,000+ 58,000 + 588,000232,000 13 14 <Prev 2 3 of 3 Next > DELL Prtscrsert Fe 7 8
Answer #1
ANTIOCH COMPANY
Financial Statements Model
Event No. Assets = Stockholders’ Equity Income Statement Statement of Cash Flow
Cash + Raw Materials + Work in Process Inventory + Finished Goods + Office Furniture + Manufacturing Equipment = Common Stock + Retained Earnings Revenue Expense = Net Income
665,000 + 70,000 + 27,000 + 58,000 + + = 588,000 + 232,000 =
1 -28000 + + + + + = + -28000 28000 = -28000 -28000 OA
2 -55000 + 55000 + + + + = + = -55000 OA
3 + -84000 + 84000 + + + = + = NA
4 -77000 + + + + + = + -77000 77000 = -77000 -77000 OA
5 -98000 + + 98000 + + + = + = -98000 OA
6 -137000 + + + + 137000 + = + = -137000 IA
7 + + + + -20000 + = + -20000 20000 = -20000 NA
8 -138000 + + + + + 138000 = + = -138000 IA
9 + + 14000 + + + -14000 = + = NA
10 -47000 + + 47000 + + + = + = -47000 OA
11 -79000 + + + + + = + -79000 79000 = -79000 -79000 OA
12 + + -258000 + 258000 + + = + = NA
13 425000 + + + + + = + 425000 425000 = 425000 425000 OA
14 + + + -174300 + + = + -174300 174300 = -174300 NA
Total 431,000 + 41,000 + 12,000 + 141,700 + 117,000 + 124,000 = 588,000 + 278,700 425,000 378,300 = 46,700 -234,000 NC